OpenStudy (mrhoola):

The life (in hours) of a computer processing unit (CPU) is modeled by a Weibull distribution with parameters beta equals 3 and delta equals 1047. Determine the mean life of the CPU (in hours)

OpenStudy (anonymous):

The Weibull distribution has the density function \[f(x;k,\lambda)=\begin{cases}\dfrac{k}{\lambda}\left(\dfrac{x}{\lambda}\right)^{k-1}\exp\left[-\left(\dfrac{x}{\lambda}\right)^{k}\right]&\text{for }x\ge0\\0&\text{for }x<0\end{cases}\] (Formula taken from the Wikipedia page - I'm not sure which is \(\beta\) and \(\delta\)) The mean will be given by \[\int_{0}^\infty xf(x)\,dx\]

OpenStudy (mrhoola):

ok . That is what I did , but the stupid wiley plus didnt accept my answer

OpenStudy (anonymous):

So is \(\beta=k\) and \(\delta=\lambda\), or the other way around?

OpenStudy (mrhoola):

thats it

OpenStudy (anonymous):

Alright, so the mean would be given by \[\int_0^\infty \dfrac{\beta}{\delta}\left(\dfrac{x}{\delta}\right)^{\beta}\exp\left[-\left(\dfrac{x}{\delta}\right)^{\beta}\right]\,dx\] I'll omit the actual values of \(\beta\) and \(\delta\) for now.

OpenStudy (anonymous):

Actually, there's a slight typo above, should be \[\int_0^\infty \dfrac{\beta}{\delta}x\left(\dfrac{x}{\delta}\right)^{\beta-1}\exp\left[-\left(\dfrac{x}{\delta}\right)^{\beta}\right]\,dx\] Substitute \(u=-\left(\dfrac{x}{\delta}\right)^\beta\), then \(du=-\dfrac{\beta}{\delta}\left(\dfrac{x}{\delta}\right)^{\beta-1}\,dx\). Now, \[\int_0^\infty \dfrac{\beta}{\delta}x\left(\dfrac{x}{\delta}\right)^{\beta-1}\exp\left[-\left(\dfrac{x}{\delta}\right)^{\beta}\right]\,dx=-\int_{0}^{-\infty} \delta\left(-u\right)^{1/\beta}\exp(u)\,du\] Substitute again, \(u=-t\) so that \(du=-dt\), then you get \[\delta\int_{0}^{\infty}t^{1/\beta}\exp(-t)\,dt\] Is this what you have?

OpenStudy (mrhoola):

no not exactly

OpenStudy (mrhoola):

I think I might have messed up on the u-sub

OpenStudy (anonymous):

When you've caught up, you'll need to know the gamma function to be able to find an exact value for the integral. \[\Gamma(n)=\int_0^\infty t^{n-1}\exp(-t)\,dt\] which almost perfectly matches what we have above. You just have to tweak the exponent a bit.